# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ III.2 Characterization of the four-photon dissipation channel
Armed with the preparation and measurement techniques described in the previous section, and after locating the frequency of the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition dressed by the pump itself [40], we investigate the time decay of population prepared in state $|4\rangle_{a}$ under pump (see Fig. 3b). Clearly, the differential signal $S_{4}$—proportional to the occupation of $|4\rangle_{a}$—decays faster as the pump amplitude increases (here referenced by the voltage $V_{p}$ applied to the mixer generating the pump). For the lowest pump amplitudes ($V_{p}\leqslant 0.1~{}$V), the decay is exponential. However, for larger pump amplitudes, we note a deviation from an exponential law, indicating that reducing the joint dynamics of the memory and buffer modes to a single-mode Lindblad dynamics governed by the dissipator $\mathcal{D}[|0\rangle_{a}\langle 4|_{a}]$ is no longer valid. We further investigate this regime with a complementary experiment in which, for the same range of pump amplitudes, we record the buffer reflection spectrum in continuous-wave (see Fig. 3c). As the pump amplitude increases, a dip in amplitude appears in the buffer line, which eventually splits into two almost separated lines. This behavior is expected at the onset of strong hybridization of the $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$ levels, which occurs when $g_{4}\sqrt{4!}\gtrsim\gamma$ with $\gamma$ the intrinsic width of the transition accounting both for homogeneous and inhomogeneous broadening.
@@
These two datasets are quantitatively reproduced (lines in Fig. 3b-c) by Lindblad master equation simulations modeling the memory and buffer interacting via the Hamiltonian (3[40]. The anharmonicity and relaxation rate of each mode are independently calibrated and accounted for in these simulations. We leave as free fit parameters the value of $g_{4}$ for each pump amplitude and the strength of the buffer drive for the whole second dataset (see Supplemental materials for details of the fit procedures and assumptions behind this model). We also adjust in our model a pure dephasing rate of $\kappa_{\phi}=2\pi\times 32$ kHz for the memory, close to the independently measured dephasing rate of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition [40]. Finally, we fit a proportionality factor between the differential signal $S_{4}$ and the occupation of state $|4\rangle_{a}$. The fitted value of $g_{4}$ is reported in Fig. 3d for each pump amplitude. It reaches $2\pi\times 185~{}$kHz for the strongest pump amplitude, corresponding to a normalized amplitude $|\epsilon(t)|=0.035$ with the convention used in the introduction. Note that the scaling of $g_{4}$ slightly deviates from the expected linear dependency against pump amplitude at room temperature, which could be explained by the saturation of the mixer generating the pump pulses.
@@
Figure 4: Four-photon loss channel. Decay rates $\Gamma_{n}$ of the first four excited states of the memory ($1\leq n\leq 4$) in presence of a pump of varying amplitude. $\Gamma_{4}$ is estimated through spectral analysis [40] of the model Lindbladian used to reproduce the datasets presented in Fig. 3b-c. Following an initial quadratic increase with respect to pump amplitude, it saturates near its maximum theoretical value at half the buffer damping rate (red dashed line). $\Gamma_{n}$ for $n\leq 3$ is directly obtained from an exponential fit of the time-decay of population prepared in $|n\rangle_{a}$ (presented as Supplemental materials). Their increase at large pump amplitude is not captured by our model.
@@
In order to quantitatively relate the largest fitted values of $g_{4}$ to four-photon dissipation rates, one needs to go beyond the simple model reduction allowed by adiabatic elimination of the buffer. We instead extract the lifetime of state $|4\rangle_{a}$ through spectral analysis of a reduced Liouvillian capturing the system dynamics [40]. We report the extracted decay rate $\Gamma_{4}$ in Fig. 4 (filled red circles). Following an initial quadratic increase of $\Gamma_{4}$ against pump amplitude—also predicted by the adiabatic elimination of the buffer—we observe a saturation of $\Gamma_{4}$ near its theoretical maximum value $\kappa_{b}/2$ set by the buffer damping rate [27]. In this regime, engineered four-photon dissipation dominates the dynamics, inducing decay of population from state $|4\rangle_{a}$ at a rate that exceeds that induced by native single-photon dissipation by a factor 10. As a sanity check, we also record, under the same conditions, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ and report the corresponding decay rates $\Gamma_{n}$ in Fig. 4. While these should nominally not be affected by the pump (see Fig. 1b), we observe a small but clear increase with pump amplitude. Such behavior was reported in similar parametrically pumped systems [23, 25], and and may stem from spurious near-resonant parametric processes [52, 53, 54, 55]. In our case, we note that a two-to-one photon exchange process between the memory and the buffer, resonant at $2\omega_{b}-\omega_{a}$, is detuned by approximately 200 MHz from the target process. We estimate that this off-resonantly driven process results in an effective pump-dependent single-photon dissipation channel for the memory, contributing marginally to the observed degradation of the low-energy levels lifetime (dissipation rate in the $0-10$ kHz range for the considered pump amplitudes). In any case, the spurious increase of $\Gamma_{n}$ for $n<4$ remains an order of magnitude lower than the recorded increase of $\Gamma_{4}$.
@@ IV.1 High-order dissipators from high-order interactions
In earlier experimental works realizing high-order dissipation channels, these were obtained from low-order interactions. Among other examples, autonomous parity recovery in a superconducting resonator was achieved by combining dispersive interactions and a set of single and two-photon drives [37], autonomous stabilization of the GKP code was achieved by combining Rabi interactions and sequences of single-photon drives [17, 19, 56], and a four-to-two photon exchange interaction—which could be used to engineer a four-photon dissipation channel—was obtained by cascading off-resonant two-to-one photon exchange interactions in a Raman configuration [57] [^6] In that work, the auxiliary mode was not designed to dissipate excitations at a sufficiently large rate to turn the interaction into an effective four-photon dissipation channel.. We argue that this approach can lead to the opening of spurious dissipation channels entailed by each low-order primitive involved in the generation of effective high-order interactions. For instance, in the Raman scheme, finite detuning of the pumps to two-to-one photon transitions can excite levels which should nominally be only virtually excited. Such excitations would directly lead to a logical flip of the encoded cat-qubit if the engineered dissipation channel were employed toward this goal.
@@ IV.2 Parameter regime
As discussed in this paper introduction, the memory and buffer phase fluctuations across the ATS should be large in order for the rate of the activated four-to-one photon exchange to be significant. However, the memory phase fluctuations cannot be chosen arbitrarily large when stabilizing four-legged cats [40]. Indeed, for a pump $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ resonant with the four-to-one photon transition, the RWA yields corrections to the exchange interaction stemming from high-order terms in the ATS sine Hamiltonian (see Eq. (2)):
@@
In this expression, :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ and $P_{a}$ and $P_{b}$ are polynomials boiling down to unit factors if the SQUID potential is truncated to leading order. These corrections yield a state-dependent four-to-one photon exchange rate which can result in instabilities when stabilizing a cat with size $\alpha\gtrsim 1/\varphi_{a}^{ZPF}$ [40, 21]. We note that this effect also arises for two-legged cats, and is one of the limiting factors to their stabilization rate  [^7] Large dissipation rates are desired for fast adiabatic control of the cat state [75] but are not needed to achieve macroscopic bit-flip times [26].. On the upside, it was recently proposed to leverage the same effect to increase the number of attractors in phase-space—thus stabilizing higher-dimension cat codes—or to stabilize more complex rotation-symmetric codes with advantageous error-correcting capabilities [21].
@@ IV.3 Spurious non-linearities and mitigation measures
In the introduction of this paper, we discussed how an idealized circuit comprising a SQUID placed in series with two resonators enabled dissipation engineering while preserving the target mode linearity. Various non-idealities can break this property and induce anharmonicity. We here outline how the observed memory anharmonicity can be traced back to stray inductance in the SQUID and give possible mitigation measures for future experiments. We refer the reader to the paper Supplemental Materials for a detailed analysis of the various non-linear mechanisms arising in our circuit.
@@
${\bf H}_{J}(t)$ $=-E_{J}\cos\big{(}\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}}\big{)}-E_{J}\cos\big{(}\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}}\big{)}$ (5)
$\qquad-E_{J_{2}}\cos\big{(}2(\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})\big{)}-E_{J_{2}}\cos\big{(}2(\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})\big{)}$
$=2E_{J}~{}\sin\big{(}\epsilon(t)\big{)}\sin(\bm{\varphi})$
$\qquad-2E_{J_{2}}~{}\cos\big{(}2\epsilon(t)\big{)}\cos(2\bm{\varphi})$
@@
We immediately see that the added contribution does not cancel at the ATS working point, even in absence of pump ($\epsilon=0$). Moreover, the vacuum phase fluctuations of each mode are effectively doubled and approach unity for this term, such that the cosine cannot be approximated by a low-order polynomial. For the fitted value $E_{J_{2}}=E_{J}/500$, we thus find a strongly non-linear perturbation of weak amplitude as sought.
@@
A known mechanism renormalizing the Josephson energy with a second harmonic stems from small stray inductances in the arms of the SQUID [47, 54]. For series inductance with inductive energy $\epsilon_{L}\gg E_{J}$ (see Fig. 2a), the amplitude of the second harmonic is $E_{J_{2}}=-E_{J}^{2}/(4\epsilon_{L})$ [40]. In our case, the observed amplitude corresponds to an inductance value of $\varphi_{0}^{2}/\epsilon_{L}=85$ pH matching our estimate for the geometric inductance of the tracks of the SQUID. Note that an alternative mechanism yielding harmonics in the Josephson potential—stemming from pinholes in the Josephson junctions oxide barrier—was recently proposed to explain reported anomalies in transmon energy level structures [69]. The amplitude of the second harmonic found in our experiment is smaller than the values reported in that work and can be entirely attributed to stray circuit inductances. This different behavior may be explained by a thicker oxide barrier in our junctions [40].
In a future iteration of the experiment, stray inductances could be significantly reduced with simple design changes. Indeed, the rather large SQUID loop in the present experiment (dimensions $\sim 50\times 50~{}\mathrm{\mu m}^{2}$) was chosen to accommodate at its center the long chain of junctions forming the shunt inductor of the ATS. However, physically placing this chain in parallel but outside of the SQUID loop would yield an equivalent circuit—simply requiring to reallocate bias fluxes between loops—while allowing to reduce the SQUID perimeter down to $\sim 10~{}\mu$m as routinely achieved in flux qubits [70]. Combined with a reduction of the junctions energy to $E_{J}=h\times 5~{}$GHz—as was originally intended in design—we estimate that the anharmonicity of the memory mode will become sufficiently low to be overcome by the four-photon dissipation.
@@
Finally, we note that while our experiment is particularly sensitive to non-linearity resulting from various circuit non-idealities (imbalance of the SQUID, non-linearity of the superinductors, harmonics in the Josephson potential), these mechanisms also exist in experiments aiming at dissipatively stabilizing other bosonic codes with similar circuits. While these experiments typically operate in a lower impedance regime in order to to suppress the anharmonicity inherited by the memory (dominant non-linear term scaling as ${\varphi_{a}^{ZPF}}^{4}\propto Z_{a}^{2}$), the rate of desired interactions decreases in the same limit. Therefore, even when targeting lower-order dissipation channels scaling more favorably with the modes impedance, the optimal parameter choice results from a tradeoff between the rate of engineered dissipation and the strength of spurious non-linearities. Our platform is an ideal testbed to investigate this tradeoff. Our experiment also highlights that the anharmonicity inherited from different mechanisms cannot be boiled down to the amplitude of a single Kerr-like term. As a result, compensating the effect of one mechanism with another as was recently realized [54] can only lead to imperfect cancellation. An intriguing exception detailed in this paper Supplemental materials is the possibility to perfectly cancel, in the RWA, the anharmonicity induced by a weak but non-zero imbalance of the SQUID with a dynamical mechanism involving off-resonant pumps. In future experiments, this scheme should prove instrumental in order not to rely on perfect SQUID symmetry as was achieved in the present experiment.

### v3

@@ III.2 Characterization of the four-photon dissipation channel
Armed with the preparation and measurement techniques described in the previous section, and after locating the frequency of the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition dressed by the pump itself (see Sec. C.2), we investigate the time decay of population prepared in state $|4\rangle_{a}$ under pump (see Fig. 3b). Clearly, the differential signal $S_{4}$—proportional to the occupation of $|4\rangle_{a}$—decays faster as the pump amplitude increases (here referenced by the voltage $V_{p}$ applied to the mixer generating the pump). For the lowest pump amplitudes ($V_{p}\leqslant 0.1$~V), the decay is exponential. However, for larger pump amplitudes, we note a deviation from an exponential law, indicating that reducing the joint dynamics of the memory and buffer modes to a single-mode Lindblad dynamics governed by the dissipator $\mathcal{D}[|0\rangle_{a}\langle 4|_{a}]$ is no longer valid. We further investigate this regime with a complementary experiment in which, for the same range of pump amplitudes, we record the buffer reflection spectrum in continuous-wave (see Fig. 3c). As the pump amplitude increases, a dip in amplitude appears in the buffer line, which eventually splits into two almost separated lines. This behavior is expected at the onset of strong hybridization of the $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$ levels, which occurs when $g_{4}\sqrt{4!}\gtrsim\gamma$ with $\gamma$ the intrinsic width of the transition accounting both for homogeneous and inhomogeneous broadening.
@@
These two datasets are quantitatively reproduced (lines in Fig. 3b-c) by Lindblad master equation simulations modeling the memory and buffer interacting via the Hamiltonian (3). The anharmonicity and relaxation rate of each mode are independently calibrated and accounted for in these simulations. We leave as free fit parameters the value of $g_{4}$ for each pump amplitude and the strength of the buffer drive for the whole second dataset (see Sec. D.2 for details of the fit procedures and assumptions behind this model). We also adjust in our model a pure dephasing rate of $\kappa_{\phi}=2\pi\times 32$ kHz for the memory, close to the independently measured dephasing rate of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition. Finally, we fit a proportionality factor between the differential signal $S_{4}$ and the occupation of state $|4\rangle_{a}$. The fitted value of $g_{4}$ is reported in Fig. 3d for each pump amplitude. It reaches $2\pi\times 185$~kHz for the strongest pump amplitude, corresponding to a normalized amplitude $|\epsilon(t)|=0.035$ with the convention used in the introduction. Note that the scaling of $g_{4}$ slightly deviates from the expected linear dependency against pump amplitude at room temperature, which could be explained by the saturation of the mixer generating the pump pulses.
@@
Figure 4: Four-photon loss channel. Decay rates $\Gamma_{n}$ of the first four excited states of the memory ($1\leq n\leq 4$) in presence of a pump of varying amplitude. $\Gamma_{4}$ is estimated through spectral analysis of the model Lindbladian used to reproduce the datasets presented in Fig. 3b-c. Following an initial quadratic increase with respect to pump amplitude, it saturates near its maximum theoretical value at half the buffer damping rate (red dashed line). $\Gamma_{n}$ for $n\leq 3$ is directly obtained from an exponential fit of the time-decay of population prepared in $|n\rangle_{a}$ (see Sec. C.3). Their increase at large pump amplitude is not captured by our model.
@@
In order to quantitatively relate the largest fitted values of $g_{4}$ to four-photon dissipation rates, one needs to go beyond the simple model reduction allowed by adiabatic elimination of the buffer. We instead extract the lifetime of state $|4\rangle_{a}$ through spectral analysis of a reduced Liouvillian capturing the system dynamics, as detailed in Sec. D.2. We report the extracted decay rate $\Gamma_{4}$ in Fig. 4 (filled red circles). Following an initial quadratic increase of $\Gamma_{4}$ against pump amplitude—also predicted by the adiabatic elimination of the buffer—we observe a saturation of $\Gamma_{4}$ near its theoretical maximum value $\kappa_{b}/2$ set by the buffer damping rate [55]. In this regime, engineered four-photon dissipation dominates the dynamics, inducing decay of population from state $|4\rangle_{a}$ at a rate that exceeds that induced by native single-photon dissipation by a factor 10. As a sanity check, we also record, under the same conditions, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ and report the corresponding decay rates $\Gamma_{n}$ in Fig. 4. While these should nominally not be affected by the pump (see Fig. 1b), we observe a small but clear increase with pump amplitude. Such behavior was reported in similar parametrically pumped systems [28, 29], and may stem from spurious near-resonant parametric processes [48, 79, 54, 7]. In our case, we note that a two-to-one photon exchange process between the memory and the buffer, resonant at $2\omega_{b}-\omega_{a}$, is detuned by approximately 200 MHz from the target process. We estimate that this off-resonantly driven process results in an effective pump-dependent single-photon dissipation channel for the memory, contributing marginally to the observed degradation of the low-energy levels lifetime (dissipation rate in the 0100-10 kHz range for the considered pump amplitudes). In any case, the spurious increase of $\Gamma_{n}$ for $n<4$ remains an order of magnitude lower than the recorded increase of $\Gamma_{4}$.
@@ IV.1 High-order dissipators from high-order interactions
In earlier experimental works realizing high-order dissipation channels, these were obtained from low-order interactions. Among other examples, autonomous parity recovery in a superconducting resonator was achieved by combining dispersive interactions and a set of single and two-photon drives [15], autonomous stabilization of the GKP code was achieved by combining Rabi interactions and sequences of single-photon drives [61, 10, 25], and a four-to-two photon exchange interaction—which could be used to engineer a four-photon dissipation channel—was obtained by cascading off-resonant two-to-one photon exchange interactions in a Raman configuration [39, 21]. We argue that this approach can lead to the opening of spurious dissipation channels entailed by each low-order primitive involved in the generation of effective high-order interactions. For instance, in the Raman scheme, finite detuning of the pumps to two-to-one photon transitions can excite levels which should nominally be only virtually excited. Such excitations would directly lead to a logical flip of the encoded cat-qubit if the engineered dissipation channel were employed toward this goal.
@@ IV.2 Parameter regime
As discussed in this paper introduction, the memory and buffer phase fluctuations across the ATS should be large in order for the rate of the activated four-to-one photon exchange to be significant. However, as detailed in Sec. A.1, the memory phase fluctuations cannot be chosen arbitrarily large when stabilizing four-legged cats. Indeed, for a pump $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ resonant with the four-to-one photon transition, the RWA yields corrections to the exchange interaction stemming from high-order terms in the ATS sine Hamiltonian (see Eq. (2)):
@@
In this expression, :_::\_: denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ and $P_{a}$ and $P_{b}$ are polynomials boiling down to unit factors if the SQUID potential is truncated to leading order. These corrections yield a state-dependent four-to-one photon exchange rate which can result in instabilities when stabilizing a cat with size $\alpha\gtrsim 1/\varphi_{a}^{ZPF}$ [59]. We note that this effect also arises for two-legged cats, and is one of the limiting factors to their stabilization rate [26]. On the upside, it was recently proposed to leverage the same effect to increase the number of attractors in phase-space—thus stabilizing higher-dimension cat codes—or to stabilize more complex rotation-symmetric codes with advantageous error-correcting capabilities [59].
@@ IV.3 Spurious non-linearities and mitigation measures
In the introduction of this paper, we discussed how an idealized circuit comprising a SQUID placed in series with two resonators enabled dissipation engineering while preserving the target mode linearity. Various non-idealities can break this property and induce anharmonicity. We here outline how the observed memory anharmonicity can be traced back to stray inductance in the SQUID and give possible mitigation measures for future experiments. We refer the reader to Sec. B for a detailed analysis of the various non-linear mechanisms arising in our circuit.
@@
${\bf H}_{J}(t)$ $=-E_{J}\cos\big(\boldsymbol{\varphi}-\frac{\Phi_{1}}{\varphi_{0}}\big)-E_{J}\cos\big(\boldsymbol{\varphi}+\frac{\Phi_{2}}{\varphi_{0}}\big)$ (5)
$\qquad-E_{J_{2}}\cos\big(2(\boldsymbol{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})\big)-E_{J_{2}}\cos\big(2(\boldsymbol{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})\big)$
$=2E_{J}~\sin\big(\epsilon(t)\big)\sin(\boldsymbol{\varphi})$
$\qquad-2E_{J_{2}}~\cos\big(2\epsilon(t)\big)\cos(2\boldsymbol{\varphi})$
@@
We immediately see that the added contribution does not cancel at the ATS working point, even in absence of pump ($\epsilon=0$). Moreover, the vacuum phase fluctuations of each mode are effectively doubled and approach unity for this term, such that the cosine cannot be approximated by a low-order polynomial. For the fitted value $E_{J_{2}}=E_{J}/500$, we thus find a strongly non-linear perturbation of weak amplitude as sought. While this term has a large impact on the memory-mode anharmonicity, raising it by roughly two orders of magnitude, it has no significant effect on the four-to-one exchange mechanism we aim to activate.
@@
A known mechanism renormalizing the Josephson energy with a second harmonic stems from small stray inductances in the arms of the SQUID [65, 54]. For series inductance with inductive energy $\epsilon_{L}\gg E_{J}$ (see Fig. 2a), the amplitude of the second harmonic is $E_{J_{2}}=-E_{J}^{2}/(4\epsilon_{L})$. In our case, the observed amplitude corresponds to an inductance value of $\varphi_{0}^{2}/\epsilon_{L}=85$ pH matching our estimate for the geometric inductance of the tracks of the SQUID (see Sec. B). Note that an alternative mechanism yielding harmonics in the Josephson potential—stemming from pinholes in the Josephson junctions oxide barrier—was recently proposed to explain reported anomalies in transmon energy level structures [83]. The amplitude of the second harmonic found in our experiment is smaller than the values reported in that work and can be entirely attributed to stray circuit inductances. This different behavior may be explained by a thicker oxide barrier in our junctions.
[…]
@@
Finally, we note that while our experiment is particularly sensitive to non-linearity resulting from various circuit non-idealities (imbalance of the SQUID, non-linearity of the superinductors, harmonics in the Josephson potential), these mechanisms also exist in experiments aiming at dissipatively stabilizing other bosonic codes with similar circuits. While these experiments typically operate in a lower impedance regime in order to to suppress the anharmonicity inherited by the memory (dominant non-linear term scaling as ${\varphi_{a}^{ZPF}}^{4}\propto Z_{a}^{2}$), the rate of desired interactions decreases in the same limit. Therefore, even when targeting lower-order dissipation channels scaling more favorably with the modes impedance, the optimal parameter choice results from a tradeoff between the rate of engineered dissipation and the strength of spurious non-linearities. Our platform is an ideal testbed to investigate this tradeoff. Our experiment also highlights that the anharmonicity inherited from different mechanisms cannot be boiled down to the amplitude of a single Kerr-like term. As a result, compensating the effect of one mechanism with another as was recently realized [54] can only lead to imperfect cancellation. An intriguing exception detailed in Sec. B is the possibility to perfectly cancel, in the RWA, the anharmonicity induced by a weak but non-zero imbalance of the SQUID with a dynamical mechanism involving off-resonant pumps. In future experiments, this scheme should prove instrumental in order not to rely on perfect SQUID symmetry as was achieved in the present experiment.